• How many grams of aluminum oxide are produced from the synthesis reaction of 30.0 g of Al ?
  • Balance the equation= _ Al + O2 -> _ Al2O3
  • Factor label( molar mass to mole , mole to mole ratio , mole to molar mass )

*This is one whole question *

Please & thank you.



1
Expert's answer
2021-05-24T07:20:00-0400

Solution:

Balanced chemical equation:

4Al + 3O2 → 2Al2O3


The molar mass of Al is 26.98 g/mol.

Therefore,

(30.0 g Al) × (1 mol Al / 26.98 g Al) = 1.112 mol Al

n(Al) = 1.112 mol


According to the equation: n(Al)/2 = n(Al2O3)

Therefore,

n(Al2O3) = n(Al)/2 = 1.112 mol / 2 = 0.556 mol Al2O3


The molar mass of Al2O3 is 101.96 g/mol.

Therefore,

(0.556 mol Al2O3) × (101.96 g Al2O3 / 1 mol Al2O3) = 56.6898 g Al2O3 = 56.7 g Al2O3


Answer: 56.7 g Al2O3